An airplane accelerates down a runway at 3.20 m/s2 for 32.8 s until is finally lifts off the ground. Determine the distance traveled before takeoff.


1
Expert's answer
2021-02-10T08:34:55-0500

Assuming that an airplane starts from rest, we have:


"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\cdot3.2\ \dfrac{m}{s^2}\cdot(32.8\ s)^2=1721.3\ m."